An artificial satellite is in a circular orbit d=730.0 km above the surface of a planet of radius r=4.55×103 km. The period of revolution of the satellite around the planet is T=2.15 hours. What is the average density of the planet? kg/m3
Answer #1

Gravitational constant = G = 6.67 x 10-11 N.m2/kg2

Mass of the planet = M

Mass of the satellite = m

Radius of the planet = Rp = 4.55 x 103 km = 4.55 x 106 m

Distance of the satellite above the surface of the planet = H = 730 km = 730 x 103 m = 0.73 x 106 m

Radius of orbit of the satellite = R

R = Rp + H

R = 4.55x106 + 0.73x106

R = 5.28 x 106 m

Orbital period of the satellite = T = 2.15 hours = 2.15 x (3600) sec = 7740 sec

Orbital speed of the satellite = V

VT = 2\piR

V(7740) = 2\pi(5.28x106)

V = 4286.204 m/s

The centripetal force for the circular motion of the satellite is provided by the gravitational force of the planet on the satellite.

\frac{mV^{2}}{R} = \frac{GMm}{R^{2}}

V^{2} = \frac{GM}{R}

M = \frac{V^{2}R}{G}

M = \frac{(4286.204)^{2}(5.28\ast 10^{6})}{6.67\ast 10^{-11}}

M = 1.454 x 1024 kg

Volume of the planet = V0

V0 = 4\piRp3/3

V0 = 4\pi(4.55x106)3/3

V0 = 3.945 x 1020 m3

Average density of the planet = \rho

M = \rhoV0

1.454x1024 = \rho(3.945x1020)

\rho = 3685.68 kg/m3

Average density of the planet = 3685.68 kg/m3